Two discharge tubs have identical material structure and the same gas is filled in them. The length of one tub e is 10 cm and that of the other tube id 20 cm. Sparking starts in both the tubes when the potential difference between the cathode and the anode is 100 V. If the pressure in the shorter tube is 1.0 mm of mercury, what is the pressure in the longer tube?

Grade:upto college level

1 Answers

Aditi Chauhan
9 years ago
Sol. V = f(Pd) v basse s = P base s d base s v base L = P base l d base l ⇒ V base s/V base l = P base s/P base l * d base s/d base l ⇒ 100/100 = 10/20 * 1mm/x ⇒ x = 1 mm / 2 = 0.5 mm